The length of a rectangular garden lawn is 8m. If the perimeter of the rectangle is 72m. what is its area?
ans:
P= 2(l+b)
72=2(8+b)
72=16+2b
b=38

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

The length of a rectangular garden lawn is 8m. If the perimeter of the rectangle is 72m. what is its area?
--------
P = 2L + 2W = 72
2*8 + 2W = 72
2W = 56
W = 28 meters
---------
Area = L*W = 8*28
Area = 224 sq meters

Answer by ikleyn(52781)

.
I always used to think that the length is the largest dimension of a rectangle (by definition), while the width is the smallest dimension.